OpenStudy (anonymous):

Write the standard equation of the circle with center (2, 3) and a diameter of 12.

jimthompson5910 (jim_thompson5910):

Hint: if the diameter is 12, then the radius is 12/2 = 6.

OpenStudy (anonymous):

(x - 2)2 + (y - 3)2 = 6?

jimthompson5910 (jim_thompson5910):

Another hint: General equation of circle: \(\Large (x-h)^2 + (y-k)^2 = r^2\)

jimthompson5910 (jim_thompson5910):

you are close, but something is off

OpenStudy (anonymous):

(x - 2)^2 + (y - 3)^2 = 6^2

jimthompson5910 (jim_thompson5910):

better, then change 6^2 into 36

OpenStudy (anonymous):

ok i got it, :)